# Towards Practical Finite Sample Bounds for Motion Planning in TAMP

### v1

@@ Abstract
When using sampling-based motion planners, such as PRMs, in configuration spaces, it is difficult to determine how many samples are required for the PRM to find a solution consistently. This is relevant in Task and Motion Planning (TAMP), where many motion planning problems must be solved in sequence. We attempt to solve this problem by proving an upper bound on the number of samples that are sufficient, with high probability, to find a solution by drawing on prior work in deterministic sampling and sample complexity theory. We also introduce a numerical algorithm to compute a tighter number of samples based on the proof of the sample complexity theorem we apply to derive our bound. Our experiments show that our numerical bounding algorithm is tight within two orders of magnitude on planar planning problems and becomes looser as the problem’s dimensionality increases. When deployed as a heuristic to schedule samples in a TAMP planner, we also observe planning time improvements in planar problems. While our experiments show much work remains to tighten our bounds, the ideas presented in this paper are a step towards a practical sample bound.
@@ 1 Introduction
Sampling-based motion planners, such as PRMs, perform well in complicated configuration spaces of moderately high dimension. It is difficult to determine how many samples a PRM needs to find a solution consistently. This problem appears when sampling-based motion planners are used as subroutines for Task and Motion Planning (TAMP), where many motion planning problems must be solved in sequence. Existing TAMP solvers address this problem by manually selecting a uniform threshold for the motion planner’s computational effort to find a solution. We seek a more principled method of estimating sample thresholds.
@@
We attempt to solve our problem by proving a bound on the number of random samples for a PRM to find a solution. First, we must choose a set of quantitative features (a problem parametrization) of the motion planning problem on which we derive our bound. Two notable problem parametrizations and sample bounds lay the foundation of our work. Tsao et al. suggest using the widths of passages in the configuration space as a parameterization [31]. They derive the number of deterministic samples required to cover the space at a specified density to yield PRMs guaranteed to have paths in these narrow passages. While their parameterization is simple, TAMP systems generally cannot afford to use deterministic sample sets. If a planner returns to a previously attempted motion planning problem, a denser set of samples must be recomputed from scratch (and previous computation is wasted) [25, 14]. Hsu et al. introduce a random sample bound based on a three-parameter problem characterization called expansiveness, which characterizes the visibility properties of individual points and connected sets in the configuration space [12]. Should we compute samples with his bound, iterative search through three parameters if the planner fails to solve the problem is a difficult task. We wish to have a simple problem parameterization, like Tsao et al., but for random sample sets like Hsu et al. that can be incrementally refined.
@@
This paper presents a sample bound that extends Tsao et al.’s radius PRM guarantee (parameterized by narrow passage width) to random samples. The core of our argument draws a connection to work in sample complexity by Blumer et al. [2] that estimates the probability of finding covering point configurations of a specific density. When Blumer et al.’s result is applied in practice, their bound significantly overestimates the number of required samples resulting in a larger PRM than what can tractably be stored. Rather than using the explicit formula stated in the theorem of Blumer et al., we present a novel numerical algorithm that searches over a tighter expression in their proof [2] to compute a tighter number of samples for a given probability of failure. We also attempt to extend our radius PRM result to the KNN PRM, but our analysis falls short of providing a well-defined bound. However, we include the intermediate analysis here to provide a starting point for a future finite-sample bound.
@@
In our experiments, we evaluate the numerical algorithm’s tightness and utility in a TAMP planner. We find that the numerical bounding algorithm is tightest in planar motion planning problems and loosens as the number of dimensions in the configuration space increases. Consequently, the bound also speeds up planning time for planar TAMP problems but shows limited improvement on higher-dimensional problems. Even so, we believe the theoretical machinery and algorithmic ideas presented in this paper are a useful step towards a practical bound.
@@ 2 Background and Problem Formulation
6:         $G_{PRM},\delta,\gamma\leftarrow\text{{RestoreProgress}}(X,progress)$triangleright $\delta$ is passage width.
@@
7:         $N_{new}\leftarrow\text{{{ComputeNewSamples}}}(\text{{vol}}(X),\delta,\gamma)$$\triangleright$ $\gamma$ is failure prob.
8:         $G_{PRM}\leftarrow\text{{sPrmConstruction}}(N_{new},X,G_{PRM})$
9:         $path\leftarrow\text{{sPrmQuery}}(s,g,G_{PRM})$
10:         $progress.\text{{update}}(X,s,g,G_{PRM})$
11:         if $path\neq\emptyset$ then
12:              $paths.\text{{append}}(path)$
13:         else
14:              $\delta^{\prime}\leftarrow\text{{FinerResolution}}(\delta)$
15:              $skeleton.\text{{UpdateMpProblem}}(X,s,g,\delta^{\prime})$
16:              $tp.\text{{QueuePush}}(skeleton)$
17:              $success\leftarrow\text{{False}}$
18:              break
19:     if $success$ then:
20:         return $paths$
21:return $\emptyset$
We aim to estimate the computational effort required to solve a constituent motion planning problem in a task skeleton proposed by a task planner. Algorithm 1 sketches out the planning routine we use in our problem setting, which we walk through in this section. We use the PRM definition as Karaman and Frazzoli [15], which is also included in the supplemental material.
The TAMP planner (Alg. 1) uses a priority queue of task plans, each of which proposes a sequence of motion planning problems to be solved by the PRM (lines 2-8). A motion planning problem consists of a configuration space $X\subset\operatorname{\mathbb{R}}^{d}$, a collision-free configuration space $X_{free}\subseteq X$, a start pose $x_{s}\in X$, and a goal pose $x_{g}\in X$, which we will denote by the tuple $(X,X_{free},x_{s},x_{g})$. We assume that $X_{free}$ is Borel measurable and can be sampled uniformly directly, so all our results and experiments will be done using $X_{free}$ directly without much reference to $X$. In practice, the measure of $X_{free}$ can be computed efficiently by Monte Carlo integration [35].
@@
Whenever sampling-based motion planners are used as a subroutine in TAMP, the planner must decide how many samples to use in each problem (line 6). Our main objective is to develop a mathematically well-founded bound on the number of samples required for a PRM to solve the problem (with high probability), and then write an efficient implementation of that bound. Our bound depends on the width $\delta$ of the passages in $X_{free}$ to find paths and the probability of failure $\gamma$ of finding a path even if one exists. Generally, a planning algorithm does not know the widths of the passages of a configuration space in advance. As a result, Alg. 1 first computes samples and PRM assuming wide paths and high failure probabilities on the initial attempt to solve each motion planning subproblem. On successive iterations of the same motion planning problem, the algorithm progressively shrinks the path width and failure probability (and thus increases the number of samples; Alg. 1, line 13). We note that in this implementation of Alg. 1, a skeleton containing an infeasible motion planning problem will always remain in the queue (but may not be given a high priority depending on the task planner). Alternatively, the minimum passage width can be set based on the robot’s control infrastructure’s capability to execute a motion plan in a narrow corridor. Should the planner fail to find a motion plan in a skeleton on an iteration using the minimum width, the planner could remove the skeleton from the queue.
@@
The only piece of notation we require is that $B_{r}^{d}(x)$ represents a $d$-dimensional ball with radius $r$ centered at point $x$. If the location of the ball does not matter (if we are only discussing its volume, etc.), we will abbreviate to $B_{r}^{d}$.
@@ 3 Random Covering Bounds for Radius PRMs
We begin by first proving a sample bound for radius PRMs. These PRMs approximate the configuration space with a set of points, where a single point represents a ball region with a radius equivalent to the connection radius. Tsao et al. observed that if a set of samples forms a sufficiently dense covering of the configuration space, the sample set will be a good approximation of the entire space. They then showed that building roadmaps from dense covering sets are guaranteed to find paths in narrow passages of a specific width. We introduce the terminology and results to state precisely what we mean by a narrow passage, a covering set of samples, and the relation between the two for the PRM guarantee. Proofs of all results in this section are included in the supplemental material.
@@
We start with the definition of narrow passages. If there exists a solution path between $x_{s}$ and $x_{g}$, then we can associate a ‘clearance’ with the path, the extent to which the robot can deviate from the path without hitting obstacles:
@@
Let $(X,X_{free},x_{0},x_{g})$ be a motion planning problem. A continuous path $\sigma:[0,1]\to X$ s.t. $\sigma(0)=x_{0}$ and $\sigma(x_{g})=1$ is $\delta$-clear if $B_{\delta}(\sigma(t))\subset X_{free}$ for all $0\leq t\leq 1$.
@@
The mathematical term for a covering set of points is called an $\alpha$-net. A common intuitive definition is first to associate each point with a ball region of radius $\alpha$ centered around the point. Should the union of the $\alpha$-balls subsume the entire configuration space (‘covers’ the configuration space), then we call the set of points an $\alpha$-net (Fig. 1). In our case, we work with an equivalent definition:
@@
Let $X\subset\operatorname{\mathbb{R}}^{n}$. A finite subset $N\subset X$ is an $\alpha$-net of $X$ if for every $x\in X$ there exists $x_{n}\in N$ such that $||x-x_{n}||_{2}\leq\alpha$.
We now state a claim proved by Tsao et al. [^1] In Tsao et al. [31], they prove a more complicated version of the claim that involves path optimality., which shows that if a set of points forms an $\alpha$-net, then a (radius) PRM with a suitable connection radius will be guaranteed find paths for paths with clearance $2\alpha$.
@@
Let $X_{free}$ be a collision-free configuration space. Suppose that $N\subset X_{free}$ is a finite set of samples that forms an $\alpha$-net of $X_{free}$, used to construct a radius PRM with connection radius $4\alpha$.
For all $x_{s},x_{g}\in X_{free}$, if there exists a solution path between $x_{s},x_{g}$ that has clearance $2\alpha$, then the resulting PRM will return a feasible path from $x_{s}$ to $x_{g}$.
A visual depiction of the proof can be found in Fig. 1c (and the full proof can be found in the supplemental material).
Tsao et al. were interested in deterministic algorithms to construct $\alpha$-nets based on layered point lattices [31]. Because TAMP planners like Alg. 1 may return to previously initiated motion planning efforts, we wish to increment previously built PRMs to save and reuse computational effort as much as possible. Deterministic sample sets are constructed with a fixed net radius $\alpha$ in mind and must be reconstructed from scratch if a finer net is required. This is why we use sample sets obtained from random sampling, which can be naturally incremented when an algorithm revisits a motion planning problem.
@@
Let $X_{free}\subset\operatorname{\mathbb{R}}^{d}$ be a collision-free configuration space, and let $\operatorname{\mathrm{P}}$ be the uniform (Lebesgue) measure on $X_{free}$ and $\gamma\in(0,1)$ be a failure probability.
Suppose $N\subset X_{free}$ is a set of $n$ i.i.d. uniform random samples. If:
$n\geq\max\left\{\frac{4}{\operatorname{\mathrm{P}}(B^{d}_{\alpha})}\log\frac{2}{\gamma},\frac{8d}{\operatorname{\mathrm{P}}(B^{d}_{\alpha})}\log\frac{13}{\operatorname{\mathrm{P}}(B^{d}_{\alpha})}\right\},$ (1)
@@
Putting Lemmas 1 and 2 together, we now have the probabilistic bound for radius-PRMs we were looking for:
@@
$n\geq\max\left\{\frac{4}{\operatorname{\mathrm{P}}(B^{d}_{\alpha})}\log\frac{2}{\gamma},\frac{8d}{\operatorname{\mathrm{P}}(B^{d}_{\alpha})}\log\frac{13}{\operatorname{\mathrm{P}}(B^{d}_{\alpha})}\right\},$ (2)
@@
$n=O\left[\frac{C_{d}}{\delta^{d}}\cdot\left(\log\frac{1}{\gamma}+d\log\frac{C_{d}}{\delta^{d}}\right)\right],\quad C_{d}=\frac{2^{d}d^{(d+1)/2}}{(2\pi e)^{d/2}}.$ (3)
We see that the number of samples varies nearly-linearly in $1/\delta$ and a dimension-dependent constant $C_{d}$, and logarithmically in $1/\gamma$. We do not find this result very surprising – a commonly held belief about PRMs is that one must sample enough to cover the entire space in the worst case, which introduces an exponential sample complexity based on dimension. Our sample complexity argument makes this idea very precise and captures a dependence on path clearance that agrees with prior work on narrow passages [11, 16].
We could use Theorem 3.1 directly by using the right side of the inequality to choose the number of samples for ComputeNewSamples in Alg. 1. In practice, this bound is quite loose. The culprit is the maximization operation in Lemma 2. The maximization’s right-hand argument shows the existence of ‘small’ $\alpha$-nets (up to constant factors) by the probabilistic method. The left-hand argument helps us estimate the probability of failure $\gamma$. The issue is that the constant on the right-hand side is too large, and so $\gamma$ must already be very small (some hand computations suggest less than $10^{-3}$ for moderate net radii) for the left-hand argument to dominate.
To find a tighter number of samples, we must bypass the maximization operation. We summarize Blumer et al.’s proof of Lemma 2 (appears as Theorem 2.1[2]) for insight, which estimates the probability that a sample set $N$ does not form an $\alpha$-net:
@@
Substituting in the sample inequality in Lemma 2 renders the last line smaller than $\gamma$, for $\gamma$ sufficiently small (Lemma A2.4 [2]).
The third line (Eq. 6) is a much tighter bound on the probability that $N$ fails to be an $\alpha$-net. We write a numerical algorithm that searches for the smallest $n=|N|$ such that the third line is below a given probability $\gamma$ (Alg. 2).
@@
1:$1>\gamma>0$, $p=\operatorname{\mathrm{P}}(B_{\alpha}^{d})$ volume of $d$-dim. $\alpha$-ball by uniform measure on c-space $X$.
@@
3:while $\left(\sum_{i=1}^{d+1}\binom{2(N_{u}+1)}{i}\right)2^{-p(N_{u}+1)/2}<\left(\sum_{i=1}^{d+1}\binom{2N_{u}}{i}\right)2^{-pN_{u}/2}<\gamma$ do
@@
7:     if $\left(\sum_{i=1}^{d+1}\binom{2(N_{mid}+1)}{i}\right)2^{-p(N_{mid}+1)/2}<\left(\sum_{i=1}^{d+1}\binom{2N_{mid}}{i}\right)2^{-pN_{mid}/2}<\gamma$ then
@@
Figure 2: Left: A comparison of the sufficient number of samples required to find an $0.5$-net with probability $1-\gamma$ as computed by Lemma 2 and numerical computation done by Algorithm 2. Most notably, the right-hand side in the maximization in Lemma 2 dominates for nearly all $\gamma$ plotted, since the dotted lines are nearly horizontal. The lines pass each other due to poor numerical precision. Right: A depiction of the search conducted by the numerical algorithm.
This algorithm exploits the fact that probability bound in Eq. 6 exhibits a monotonically increasing and then decreasing behavior as a function of the number of samples. The algorithm uses a doubling scheme to first find a conservative upper bound on the sufficient number of samples on the decreasing side of the global maximum. A binary search narrows down the correct sample count (also on the decreasing side of the global maximum, Fig. 2b). A proof of the algorithm’s correctness may be found in the supplemental material. In practice, we find that using our method is substantially tighter than using the result directly from Lemma 2, and allows us to express the scaling relationship of the number of samples with the failure probability $\gamma$ (Fig. 2b). We do find some looseness in the numerical bound, which is why some of the lines cross each other. This occurs because our implementation experiences some numerical issues when converting the large combinatorial sum in Eq. 6 and small $\gamma$ to log-space for comparison.
The runtime of the algorithm is output-sensitive. Suppose $n$ is the output number of samples. The doubling computation will double $N_{u}$ until $N_{u}\geq n$. Then, the binary search will search across the integers $[N_{u}]$. On every iteration of the doubling and search phases of the algorithm, we sum $d$ binomial coefficient computations. A binomial coefficient computation $\binom{m}{k}$ runs in $O(m)$. Putting everything together, we can express the runtime as the summation $2\sum_{i=0}^{\left\lceil\log n\right\rceil}d2^{i}=O(dn)$.

### v3

@@ Abstract
When using sampling-based motion planners such as PRMs, it is difficult to determine how many samples are required for the PRM to find a solution consistently. This is particularly relevant in Task and Motion Planning (TAMP), where many motion planning problems must be solved sequentially. We attempt to address this problem by proving an upper bound on the number of samples that are sufficient, with high probability, for a radius PRM to find a feasible solution, drawing on prior work in deterministic sampling and sample complexity theory. We also introduce a numerical algorithm that refines the bound based on the proofs of the sample complexity results we leverage. Our experiments show that our numerical algorithm is tight up to two to three orders of magnitude on planar problems for radius PRMs but becomes looser as the problem’s dimensionality increases. The numerical algorithm is empirically more useful as a heuristic for estimating the number of samples needed for a KNN PRM in low dimensions. When deployed to schedule samples for a KNN PRM in a TAMP planner, we also observe improvements in planning time in planar problems. While our experiments show that much work remains to tighten our bounds, the ideas presented in this paper are a step towards a practical sample bound.
@@ 1 Introduction
Sampling-based motion planners, such as PRMs, perform well in configuration spaces of moderately high dimension. It is difficult, however, to determine how many samples a PRM needs to find a solution consistently. This problem is critical when sampling-based motion planners are used for Task and Motion Planning (TAMP), where many motion planning problems must be solved in sequence. Existing TAMP solvers address this problem by manually selecting a uniform threshold on the motion planner’s computational effort on a problem [10, 9]. We seek a more principled method of estimating sample thresholds.
@@
We attempt to address our problem by proving a bound on the number of random samples for a PRM to find a solution with high probability. First, we must choose a parametrization of the motion planning problem on which we derive our bound. Our work builds on two problem parametrizations and sample bounds. Tsao et al. suggest using the widths of passages in the configuration space as a parameterization [31]. They derive the size of a deterministic sample set that covers the space at a specified density to yield PRMs guaranteed to have paths in these narrow passages. While their parameterization is simple, TAMP systems generally cannot afford to use deterministic sample sets. Since the actual widths of passages are unknown a priori, many planners return to a previously attempted motion planning problem to refine the graph when a motion plan has not been found on a previous iteration. The refined graph must consist of a denser set of deterministic samples that must be recomputed from scratch, wasting previous computation [27, 15]. Hsu et al. introduce a bound for random samples, which can be incrementally refined. Their bound uses three parameters that describe the configuration space’s ‘expansiveness,’ or visibility properties points and connected sets [13]. This complex parameterization complicates the choice of a procedure to refine the parameters when the planner reattempts a problem. We wish to have a simple problem parameterization, such as the one provided by Tsao et al., but for random sample sets that can be incrementally constructed in the manner of Hsu et al.
@@
This paper presents a sample bound that extends Tsao et al.’s radius PRM guarantee to random samples. The core of our argument draws a connection to work in sample complexity by Blumer et al. [2] that estimates the probability of finding covering point configurations of a specific density. When Blumer et al.’s result is applied in practice, their bound significantly overestimates the number of sufficient samples. Rather than using the explicit formula stated in the theorem of Blumer et al., we present a numerical algorithm that searches over a tighter expression in their proof [2] to compute a less conservative estimate of the number of samples. We also extend our radius PRM result to the KNN PRM. Our analysis falls short of a non-vacuous bound on the number of sufficient samples, but we include the analysis here for a future finite-sample bound.
@@
Our experiments evaluate the utility of the numerical algorithm in a TAMP planner. In radius PRMs, we find that the numerical bounding algorithm is tightest in planar motion planning problems and loosens as the number of dimensions in the configuration space increases. However, the same numerical algorithm appears to be more effective as a heuristic to guide the sampling effort of a KNN PRM than as a guarantee for the radius PRM itself in low dimensions. Consequently, the bound also speeds up planning time for planar TAMP problems but shows limited improvement on higher-dimensional problems. Even so, we believe the theoretical machinery and algorithmic ideas presented in this paper are a step towards a practical bound.
@@ 2 Background and Problem Formulation
We aim to estimate the sampling effort required to solve a constituent motion planning problem in a task skeleton proposed by a task planner. Alg. 1 sketches out a general TAMP planning strategy we use in our problem setting, which we walk through in this section. We use the same PRM definition as Karaman and Frazzoli [16], which we also include in the supplemental material for completeness.
The TAMP planner (Alg. 1) uses a priority queue of task plans or skeletons, each of which proposes a sequence of motion planning problems to be solved by the PRM (lines 5-11). If solutions to the motion planning problems are found, they can be combined with the skeleton for a complete solution to the planning problem. A motion planning problem consists of a configuration space $X\subset\operatorname{\mathbb{R}}^{d}$, a collision-free configuration space $X_{free}\subseteq X$, a start pose $x_{s}\in X$, and a goal pose $x_{g}\in X$, which we will denote by the tuple $(X,X_{free},x_{s},x_{g})$. We assume that $X_{free}$ is Borel measurable and can be sampled uniformly directly, so all our theoretical results and experiments will be done using $X_{free}$ directly without much reference to $X$. In practice, the measure of $X_{free}$ can be well-approximated efficiently by Monte Carlo integration [3, 34].
@@
6:         $G_{PRM},\delta,\gamma\leftarrow\text{{RestoreProgress}}(progress)$triangleright $\delta$ is passage width.
@@
7:         $N_{new}\leftarrow\text{{{ComputeNewSamples}}}(\text{{vol}}(X_{free}),\delta,\gamma)$$\triangleright$ $\gamma$ is failure prob.
8:         $G_{PRM}\leftarrow\text{{sPrmConstruction}}(N_{new},X_{free},G_{PRM})$
9:         $path\leftarrow\text{{sPrmQuery}}(x_{s},x_{g},G_{PRM})$
10:         $progress.\text{{update}}(X_{free},x_{s},x_{g},G_{PRM})$
11:         if $path\neq\emptyset$ then $paths.\text{{append}}(path)$ else
12:             $\delta^{\prime},\gamma^{\prime}\leftarrow\text{{AdjustWidthAndFailure}}(\delta)$
13:             $skeleton.\text{{UpdateMpProblem}}(X_{free},x_{s},x_{g},\delta^{\prime},\gamma^{\prime})$
14:             $tp.\text{{QueuePush}}(skeleton)$
15:             $success\leftarrow\text{{False}}$
16:             break
17:     if $success$ then:
18:         return $paths$
@@
Whenever sampling-based motion planners are used as a subroutine in TAMP, the planner must decide how many samples to use in each problem (line 7). Our main objective is to develop a mathematically well-founded bound on the number of samples required for a PRM to solve the problem (with high probability) and then write an efficient implementation of that bound. Our bound depends on the width $\delta$ of the passages in $X_{free}$ to find paths and the probability $\gamma$ of failing to find a path even if one exists. As a result, Alg. 1 first computes samples by assuming wide paths on the initial attempt to solve each motion planning subproblem with some initial probability. On successive iterations of the same motion planning problem, the algorithm progressively shrinks the path width and decreases the probability of failing to find a path (thus increasing the number of samples; line 12). We leave the problem of using the local failure probabilities to compute a global failure probability of the whole planning algorithm as an interesting direction for future work. In this implementation of Alg. 1, a skeleton containing an infeasible motion planning problem will always remain in the queue (but may not be given a high priority depending on the task planner). Alternatively, a minimum feasibility threshold can be set using a minimum passage width based on the known limitations of the robot’s control infrastructure to follow a trajectory in a narrow corridor. Should the planner fail to find a motion plan in a skeleton on an iteration using the minimum width, the planner could remove the skeleton from the queue.
@@
The only notation we require is that $B_{r}^{d}(x)$ represents a $d$-dimensional ball with radius $r$ centered at point $x$. If the location of the ball does not matter (if we are only discussing its volume, etc.), we will abbreviate to $B_{r}^{d}$.
@@ 3 Random Covering Bounds for Radius PRMs
We begin by first proving a sample bound for radius PRMs. These PRMs approximate the configuration space with a set of points, where a single point represents a ball region with a radius equivalent to the connection radius. Tsao et al. observed that if a set of samples forms a sufficiently dense covering of the configuration space, the sample set will be a good approximation of the entire space. They then showed that roadmaps built from dense covering sets are guaranteed to find paths in narrow passages of a specific width. In this section, we introduce the terminology to define precisely what we mean by a narrow passage, a covering set of samples, and the relation between the two for the PRM guarantee. Proofs of all results in this section are in the supplemental material. [^1] Supplemental material are made available at https://sageshoyu.github.io/
assets/wafr-2024-supp.pdf
@@
We start with the definition of narrow passages. If there exists a collision-free solution from $x_{s}$ to $x_{g}$, then we can associate a ‘clearance’ with the path, the extent to which the robot can deviate from the path without hitting obstacles:
@@
Let $(X,X_{free},x_{0},x_{g})$ be a motion planning problem. A continuous path $\sigma:[0,1]\to X$ s.t. $\sigma(0)=x_{0}$ and $\sigma(1)=x_{g}$ is $\delta$-clear if $B_{\delta}(\sigma(t))\subset X_{free}$ for all $0\leq t\leq 1$.
@@
The mathematical term for a covering set of points is called an $\alpha$-net. A common intuitive definition is first to associate each point with a ball region of radius $\alpha$ centered around the point. Should the union of the $\alpha$-balls subsume the entire configuration space (‘covers’ the configuration space), then we call the set of points an $\alpha$-net (Fig. 1). We work with an equivalent definition:
@@
A finite subset $N\subset X_{free}$ is an $\alpha$-net of $X_{free}$ if for every $x\in X_{free}$ there exists $x_{n}\in N$ such that $||x-x_{n}||_{2}\leq\alpha$.
We now state a claim proved by Tsao et al. A visual depiction of the proof can be found in Fig. 1c (and the full proof can be found in the supplemental material).
@@
Suppose that $N\subset X_{free}$ is a finite set of samples that forms an $\alpha$-net of $X_{free}$, used to construct a radius PRM with connection radius $4\alpha$. For all $x_{s},x_{g}\in X_{free}$, if there exists a solution path between $x_{s},x_{g}$ that has clearance $2\alpha$, then the resulting PRM will return a feasible path from $x_{s}$ to $x_{g}$.
Tsao et al. were interested in deterministic algorithms to construct $\alpha$-nets based on layered point lattices [31]. However, their sample sets are constructed using a fixed net radius $\alpha$ and must be reconstructed from scratch if a finer net is required. Because TAMP planners like Alg. 1 may return to previously initiated motion planning efforts, we wish to increment previously built PRMs to save and reuse computational effort as much as possible. In contrast to deterministic sets, collections of samples obtained from random sampling can be naturally incremented when an algorithm revisits a motion planning problem.
@@
Let $X_{free}\subset\operatorname{\mathbb{R}}^{d}$ be a collision-free configuration space, and let $\operatorname{\mathrm{P}}$ be the uniform (Lebesgue) measure on $X_{free}$ and $\gamma\in(0,1)$ be a failure probability. Suppose $N\subset X_{free}$ is a set of $n$ i.i.d. uniform random samples. If:
$n\geq\max\left\{\frac{4}{\operatorname{\mathrm{P}}(B^{d}_{\alpha})}\log_{2}\frac{2}{\gamma},\frac{8d}{\operatorname{\mathrm{P}}(B^{d}_{\alpha})}\log_{2}\frac{13}{\operatorname{\mathrm{P}}(B^{d}_{\alpha})}\right\},$ (1)
@@
Putting Lemmas 1 and 2 together, we now have a bound for radius PRMs.
@@
$n\geq\max\left\{\frac{4}{\operatorname{\mathrm{P}}(B^{d}_{\alpha})}\log_{2}\frac{2}{\gamma},\frac{8d}{\operatorname{\mathrm{P}}(B^{d}_{\alpha})}\log_{2}\frac{13}{\operatorname{\mathrm{P}}(B^{d}_{\alpha})}\right\},$ (2)
@@
$n=O\left[\frac{C_{d}}{\delta^{d}}\cdot\left(\log_{2}\frac{1}{\gamma}+d\log_{2}\frac{C_{d}}{\delta^{d}}\right)\right],\quad C_{d}=\frac{2^{d}d^{(d+1)/2}}{(2\pi e)^{d/2}}.$ (3)
The number of samples varies nearly linearly in $(1/\delta)^{d}$ and a dimension-dependent constant $C_{d}$, and logarithmically in $1/\gamma$. This result is not surprising; since one must sample enough to cover the entire space in the worst case, the sample complexity must be exponential relative to the dimension of the configuration space. Our sample complexity argument makes this idea precise and agrees with prior work on narrow passages [12, 17].
We could use Theorem 3.1 directly by using the right side of the inequality to choose the number of samples for ComputeNewSamples in Alg. 1. In practice, this bound is quite loose. The culprit is the maximization operation in Eq. 1 in Lemma 2. The maximization’s right-hand argument shows the existence of ‘small’ $\alpha$-nets (up to constant factors) by the probabilistic method. The left-hand argument helps us estimate the probability of failure $\gamma$. The constant on the right-hand side is relatively large, so $\gamma$ must be small (some hand computations suggest less than $10^{-3}$ for moderate net radii) for the left-hand argument to dominate. To find a tighter number of samples, we must bypass the maximization operation. We summarize Blumer et al.’s proof of Lemma 2 for insight, which estimates the probability that a sample set $N$ does not form an $\alpha$-net:
@@
Substituting in the sample inequality in Lemma 2 renders the last line smaller than $\gamma$, for $\gamma$ sufficiently small (Blumer et al. Lemma A2.4). The third line (Eq. 6) is a much tighter bound on the probability that $N$ fails to be an $\alpha$-net. We can now write a numerical algorithm that searches for the smallest $n=|N|$ such that the third line is below a given probability $\gamma$ (Alg. 2).
@@
1:$1>\gamma>0$, $p=\operatorname{\mathrm{P}}(B_{\alpha}^{d})$ volume of $d$-dim. $\alpha$-ball by uniform measure on c-space $X$, $\text{{SearchExp}}(n)=\left(\sum_{i=1}^{d+1}\binom{2n}{i}\right)2^{\frac{-pn}{2}}$ the tighter exp. from Eq. 6.
@@
3:while $\neg(\text{{SearchExp}}(N_{u}+1)<\text{{SearchExp}}(N_{u})<\gamma)$ do
@@
7:     if $\text{{SearchExp}}(N_{mid}+1)<\text{{SearchExp}}(N_{mid})<\gamma$ then
@@
Figure 2: Left: A comparison of the sufficient number of samples required to find an $0.5$-net in a unit cube with probability $1-\gamma$ as computed by Lemma 2 and numerical computation done by Algorithm 2. The right-hand side in the maximization in Lemma 2 dominates for nearly all $\gamma$ plotted, since the dotted lines are nearly horizontal. Right: A depiction of the search conducted by the numerical algorithm.
This algorithm exploits the fact that probability bound in Eq. 6 exhibits a monotonically increasing and then decreasing behavior as a function of the number of samples. The algorithm uses a doubling scheme to first find a conservative upper bound on the sufficient number of samples on the decreasing side of the global maximum. A binary search narrows down the correct sample count (also on the decreasing side of the global maximum, Fig. 2). A proof of the algorithm’s correctness can be found in the supplemental material. In practice, we find that using our method results in smaller estimates of sufficient sample counts than using the number of samples predicted directly by Lemma 2, and allows us to express the scaling relationship of the number of samples with the failure probability $\gamma$ (Fig. 2). All comparisons using Eq. 6 are done in log-space to improve the numerics of the procedure. Our implementation used double-precision floats; arbitrary-precision number representations yielded meager improvements to the computed number of samples.
The runtime of the algorithm is output-sensitive. Suppose $n$ is the output number of samples. The doubling computation will grow $N_{u}$ until $N_{u}\geq n$. Then, the binary search will search across the integers $[N_{u}]$. On every iteration of the doubling and search phases of the algorithm, we sum $d$ binomial coefficient computations. A binomial coefficient computation $\binom{m}{k}$ runs in $O(m)$. Putting everything together, we can express the runtime as the summation $O(2\sum_{i=0}^{\left\lceil\log_{2}n\right\rceil}d2^{i})=O(dn)$.